Find the area common to the parabolas y2 = 4x and x2 = 4y.


1
Expert's answer
2021-05-12T05:05:39-0400

Find the area common to the parabolas "y^2=4x" and "x^2=4y" .

Solution:



Points of interception:

"\begin{cases}
 y=\pm2\sqrt x \\
 y=\frac14x^2
\end{cases}"

"\begin{cases}
 x=0 \\
 y=0
\end{cases}" and "\begin{cases}
 x=4 \\
 y=4
\end{cases}" .

Area:

"S=\displaystyle\int_0^4 (2\sqrt x-\frac14x^2)dx=" "(2\cdot \frac23\cdot x^\frac32-\frac14\cdot\frac13\cdot x^3)|_0^4=\frac{2^5}{3}-\frac{4^2}{3}=\frac{16}{3}" .

Answer: "S=\frac{16}{3}" .